For a hydrogen-like atom, give all possible spectral terms for n = 2 and n = 3.
1
Expert's answer
2019-03-11T11:04:56-0400

The Rydberg formula for  wavelengths of spectral lines is as follows


"\frac{1}{\lambda}=RZ^2\left(\frac{1}{n^2}-\frac{1}{n'^2}\right)"

Here n is the lower energy level and n' is the upper energy level.

In the case n'=2 we get Balmer series:


"n=3,4,5,\ldots, \infty"

In the case n'=3 we get Paschen series:

"n=4,5,6,\ldots, \infty"